Sally is ten years younger than Mary. Ten years ago, Sally was half as old as Mary was then. Find their present ages.
Answer by ewatrrr(24785)


Hi,
*Note: Sally always has been (and will be) 10 years younger than Mary
Let x represent Marys age. then Sally's age would be(x-10)
Ten years ago***
(x-10) = x/2
2(x-10)= x
2x - 20 = x
x = 20 Mary's age 10 year ago(Sally was 10 then)
Presently Mary is now 30 and Sally is 20
checking our answer
10 = 20/2